# The infinitesimal structure of manifolds with non-continuous Riemannian metrics

### v1

@@ 4.2 A first example of non-infinitesimal Hilbertianity
Now note that for $\mathcal{L}^{1}$-almost all $x\in\mathbb{R}$, it holds that $x\in O$ or $x$ is a Lebesgue point of $\{\theta=2\}$. Write $N^{(2)}\subset\mathbb{R}$ for the null set of points satisfy neither of that and define the $\mathcal{L}^{d}$-null set $N_{2}:=N^{(2)}\times\mathbb{R}^{d-1}\cup(\partial(0,1)^{d})$. We now have characterised the behaviour of the metric speed for almost all points in the manifold $M$.
Consider the smooth functions $f_{1},f_{2},f_{3},f_{4}:M=\mathbb{R}^{d}\to\mathbb{R}$ given by
@@
Using the results on smooth Riemannian manifolds, we get that in $O\times(0,1)^{d-1}\cup([0,1]^{d})^{c}$ the minimal weak upper gradient of all functions equal the Euclidean norm of their gradients (see Proposition 2.8).
Now again pick a point $y\in(0,1)^{d}$ such that $x_{1}(y)$ is a Lebesgue point of $\{\theta=2\}$. Pick a vector $0\neq u\in\mathbb{R}^{d}$ and compute
@@
[…]
@@
$\frac{|\partial_{u}f_{1}|}{\alpha(u)}=\frac{|u_{1}|}{|u_{1}|+|u^{\prime}|_{euc}}\leq\frac{|u_{1}|}{2|u_{1}|+(|u^{\prime}|_{euc}-|u_{1}|)}=\frac{1}{2},$ (4.29)
@@
Here, in (4.29), we used that $|u_{1}|<|u^{\prime}|_{euc}$ which implies that $|u_{1}|>0$ or $|u^{\prime}|_{euc}-|u_{1}|>0$. In (4.30), (4.31), and (4.28), one can see that $u=u_{(2)}:=e_{2}$, $u=u_{(3)}:=e_{1}+e_{2}$, and $u=u_{(4)}:=e_{1}-e_{2}$ yield equality. We now define the functions $G_{1},G_{2},G_{3},G_{4}:\mathbb{R}^{d}\to[0,\infty]$ via
@@
Let $\gamma:[0,1]\to M$ be absolutely continuous and transversal to a $\mathcal{L}^{d}$-null set $N\supset N_{g}\cup N_{2}$. As $f_{i}\in C^{1}$, we may apply the chain rule. We have that for almost every $t\in[0,1]$, it holds that $\gamma_{t}\in([0,1]^{d})^{c}\cup(O\times(0,1)^{d-1})$ or $\gamma_{t}\in((0,1)\setminus O)\times(0,1)^{d-1}$. In the first case, we get that
$\Big{|}\frac{{\rm d}}{{\rm d}s}f_{i}\circ\gamma\Big{|}_{s=t}\Big{|}=|\partial_{\dot{\gamma}_{t}}f_{i}(\gamma_{t})|=|\langle\nabla_{g}f_{i},\dot{\gamma}_{t}\rangle_{g}|\leq|\nabla_{g}f_{i}|_{g}|\dot{\gamma}_{t}|_{g}=G_{i}(\gamma_{t})|\dot{\gamma}_{t}|.$ (4.34)
@@
We next want to see that these are indeed the minimal weak upper gradients. For that we need the following result from [28, Lemma 4.23].
@@
$F_{v,\delta,x}:t\to\frac{1}{\mathcal{L}^{d}(B_{\delta}(x))}\int_{B_{\delta}(x)}k(y+tv)\,{\rm d}\mathcal{L}^{n}(y),\quad\forall t,\;|t|<1.$
@@
$F_{u_{(i)},\delta,x}:t\to\frac{1}{\mathcal{L}^{d}(B_{\delta}(x))}\int_{B_{\delta}(x)}|Df_{i}|_{w}(y+tu_{(i)})\,{\rm d}\mathcal{L}^{n}(y),\ \mathrm{and}$
$G_{u_{(i)},\delta,x}:t\to\frac{1}{\mathcal{L}^{d}(B_{\delta}(x))}\int_{B_{\delta}(x)}G_{i}(y+tu_{(i)})\,{\rm d}\mathcal{L}^{n}(y),$
@@
$\frac{1}{\mathcal{L}^{d}(B_{\delta}(x))}\int_{B_{\delta}(x)}||Df_{i}|_{w}(y)-|Df_{i}|_{w}(x)|\,{\rm d}\mathcal{L}^{n}(y)\leq\frac{\varepsilon}{8},\ \mathrm{and}$ $\frac{1}{\mathcal{L}^{d}(B_{\delta}(x))}\int_{B_{\delta}(x)}|G_{i}(y)-G_{i}(x)|\,{\rm d}\mathcal{L}^{n}(y)\leq\frac{\varepsilon}{8}.$ (4.37)
@@
$\frac{1}{2\tau}\frac{1}{\mathcal{L}^{d}(B_{\delta}(x))}\int_{-\tau}^{\tau}\Bigg{|}\int_{B_{\delta}(x)}|Df_{i}|_{w}(y+tu_{(i)})\,{\rm d}\mathcal{L}^{n}(y)-\int_{B_{\delta}(x)}|Df_{i}|_{w}(y)\,{\rm d}\mathcal{L}^{n}(y)\Bigg{|}\,{\rm d}t\leq\frac{\varepsilon}{8},\ \mathrm{and}$
$\frac{1}{2\tau}\frac{1}{\mathcal{L}^{d}(B_{\delta}(x))}\int_{-\tau}^{\tau}\Bigg{|}\int_{B_{\delta}(x)}G_{i}(y+tu_{(i)})\,{\rm d}\mathcal{L}^{n}(y)-\int_{B_{\delta}(x)}G_{i}(y)\,{\rm d}\mathcal{L}^{n}(y)\Bigg{|}\,{\rm d}t\leq\frac{\varepsilon}{8}.$ (4.38)
@@
$\frac{1}{2\tau}\frac{1}{\mathcal{L}^{d}(B_{\delta}(x))}\int_{-\tau}^{\tau}\int_{B_{\delta}(x)}G_{i}(y+tu_{(i)})-|Df_{i}|_{w}(y+tu_{(i)})\,{\rm d}\mathcal{L}^{n}(y){\rm d}t\geq\varepsilon-\frac{4\varepsilon}{8}=\frac{\varepsilon}{2}.$ (4.39)
@@
$=\frac{1}{\mathcal{L}^{d}(B_{\delta}(x))}\int_{-\tau}^{\tau}\int_{B_{\delta}(x)}(G_{i}(y+tu_{(i)})-|Df_{i}|_{w}(y+tu_{(i)}))|\dot{\gamma}^{y,i}_{t}|\,{\rm d}\mathcal{L}^{n}(y){\rm d}t\geq\frac{\tau\varepsilon}{\sqrt{2}}.$ (4.43)
Now we observe that for all points $y\in\mathbb{R}^{d}$, $t\in[0,1]$ such that $z:=y+tu_{(i)}\in(O\times(0,1)^{d-1})\cup([0,1]^{d})^{c}$, we have that $g\equiv Id$ a neighbourhood of $z$ and $\nabla_{g}f_{i}(z)=u_{(i)}$. Then by definition we have that $G_{i}(z)|\dot{\gamma}_{t}^{y,i}|=G_{i}|u_{(i)}|_{euc}=\langle\nabla_{euc}f_{i},u_{(i)}\rangle_{euc}=\partial_{u_{(i)}}f_{i}.$ Similarly, we have that for all points $y\in\mathbb{R}^{d}$, $t\in[0,1]$ such that $z:=y+tu_{(i)}\in(N_{2}\cup N_{g})^{c}\setminus(O\times(0,1)^{d-1}\cup([0,1]^{d})^{c})$ that $\partial_{u_{(i)}}f_{i}(z)=G_{i}(z)|\dot{\gamma}_{t}^{y,i}|$ (see (4.35) and the equality cases of (4.25), (4.30), (4.31), and (4.32)).
@@
Pick a function $\phi\in C_{c}^{\infty}(\mathbb{R}^{d})$ such that $\phi\equiv 1$ on $[-1,2]^{d}$. Consider the functions $\tilde{f}_{i}:=\phi f_{i}$ for $1\leq i\leq 4$. Then by the previous observations, we have that
@@
Moreover, we have that $\tilde{f}_{3}=\tilde{f}_{1}+\tilde{f}_{2}$ and $\tilde{f}_{4}=\tilde{f}_{1}-\tilde{f}_{2}$ and all four functions are in $W^{1,2}(\mathbb{R}^{d})$. But as $\mathcal{L}^{d}((0,1)^{d}\setminus O\times(0,1)^{d-1})\geq 1-2\kappa>0$ and the fact that $\mathrm{vol}_{g}$ and $\mathcal{L}^{d}$ are $2^{\frac{d}{2}}$-equivalent, we get that
@@ 4.3 An example of a Sobolev Riemannian metric, neither quasi-Riemannian nor infinitesimally Hilbertian
Fix a $d\geq 3$ and a $p\in(1,d-1)$. We will now refine the previous examples and construct a $d$-manifold $M$ and a metric $g$ such that $g,g^{-1}\in L^{\infty}_{\rm loc}$, $g\in W^{1,p}_{\rm loc}$ such that $(M,\mathsf{d}_{g},g)$ is not quasi-Riemannian and such that the metric measure space $(M,\mathsf{d}_{g},\mathrm{vol}_{g})$ is not infinitesimally Hilbertian.
Define $\omega_{d-1}:=\mathcal{L}^{d-1}(B^{\mathbb{R}^{d-1}}_{1}(0))$. Pick a decreasing sequence $(R_{m})_{m\in\mathbb{N}}\subset(0,1)$ such that
@@
Fix a function $\eta\in C^{\infty}(\mathbb{R}^{d-1},[1,2])$ that satisfies $\eta|_{B_{1/2}(0)}\equiv 1$, $\eta|_{\mathbb{R}^{d-1}\setminus B_{3/4}(0)}\equiv 2$. We inductively define the finite sets $D_{m}\subset(0,1)^{d-1}$, open sets $O_{m}\subset(-1,2)^{d-1}$, a decreasing null sequence $(r_{m})_{m\in\mathbb{N}}\subset(0,1)$ and functions $\theta_{i}\in C^{\infty}(\mathbb{R}^{d-1})$ as follows:
$D_{0}:=O_{0}:=\emptyset,\ \theta_{0}\equiv 2\in C^{\infty}(\mathbb{R}^{d-1}),R_{0}:=\frac{1}{2},r_{0}:=\frac{1}{4}.$
@@
Moreover, note that from $D_{m}\subset\frac{1}{2^{m+1}}\mathbb{Z}^{d-1}$, together with (4.49) and (4.53), we have that for $x,x^{\prime}\in D_{m}$, $x\neq x^{\prime}$, it holds
@@
$\psi_{m}(y):=\left\{\begin{array}[]{ll}2&\mathrm{if}\ y\in O_{m}^{c},\\ \eta\big{(}\frac{5(y-x)}{r_{m}}\big{)}&\mathrm{if}\ y\in B_{r_{m}}(x)\ \mathrm{for\ some\ }x\in D_{m}.\end{array}\right.$ (4.59)
@@
$\lim_{r\to 0}\frac{\sum_{m\in\mathbb{N}}r_{m}\#\{x\in D_{m}:\overline{B_{r_{m}/5}(x)}\cap\overline{B_{r}(y)}\}}{r}=0\ \mathrm{and}$ (4.65)
$\lim_{r\to 0}\frac{\sum_{m\in\mathbb{N}}r^{d-1}_{m}\#\{x\in D_{m}:\overline{B_{r_{m}/5}(x)}\cap\overline{B_{r}(y)}\}}{r^{d-1}}=0.$ (4.66)
@@
$\zeta_{m}(y,r):=\#\{x\in D_{m}:\overline{B_{r_{m}/5}(x)}\cap\overline{B_{r}(y)}\}.$ (4.67)
@@
$\zeta_{m}(y,r)$$\leq\#\big{\{}\frac{1}{2^{m+1}}\mathbb{Z}^{d-1}\cap\overline{B_{r+\frac{r_{m}}{5}}(y)}\big{\}}\leq C(d)2^{(d-1)(m+1)}\frac{5^{d-1}}{4^{d-1}}r^{d-1},$
@@
[…]
Now the fact that $r_{m}>0$ together with (4.51) imply that the right-hand-side of (4.69) converges to 00 as $r\to 0$, which implies (4.66). Moreover,
[…]

### v3

@@ 4.2 An introductory example of non-infinitesimal Hilbertianity
This completes the description of the metric speed of curves through points $y$ that are Lebesgue points of $\{g=2\}$ depending on the derivative $u$. This can be seen as the explicit expression of what is called the “derivative” of the distance function in [10].
For any point $y\in(0,1)^{d}$ such that $x_{1}(y)\in O$, we have that $g\equiv\mathrm{Id}$ in a neighbourhood of $y$. Let $\gamma:[0,1]\to M$ be an absolutely continuous curve through $y$ and fix $t\in[0,1]$ such that $\gamma_{t}=y$. Suppose that $\gamma$ is differentiable at time $t$. The metric speed $|\dot{\gamma}_{t}|$ of $\gamma$ at time $t$ depends only on the values of $\gamma$ at an arbitrarily small neighbourhood $I\subset[0,1]$ of $t$. The continuity of $\gamma$ allows to assume $\gamma(I)\subset O\times\mathbb{R}^{d-1}$. As $g$ is constant, and hence smooth, in $O\times\mathbb{R}^{d-1}$, the results on smooth Riemannian metrics (see also [4, Proposition 4.10]) show that in this case $|\dot{\gamma}_{t}|=|\dot{\gamma}_{t}|_{g(y)}=|\dot{\gamma}_{t}|_{euc}$.
Similarly, for every $y\in([0,1]^{d})^{c}$, we have that $g\equiv 2\mathrm{Id}$ in a neighbourhood of $y$. Hence, for every absolutely continuous curve $\gamma$ through $y$ with $t$ as above, we have that $|\dot{\gamma}_{t}|=|\dot{\gamma}_{t}|_{g(y)}=\sqrt{2}|\dot{\gamma}_{t}|_{euc}$.
Now note that for $\mathcal{L}^{1}$-almost all $x\in\mathbb{R}$, it holds that $x\in O$ or $x$ is a Lebesgue point of $\{\theta=2\}$. Write $N^{(2)}\subset\mathbb{R}$ for the null set of points satisfying neither of that and define the $\mathcal{L}^{d}$-null set $N_{2}:=N^{(2)}\times\mathbb{R}^{d-1}\cup(\partial(0,1)^{d})$. In the spirit of Proposition 3.3, we have now for almost all points $y$, more precisely all points $y\in\mathbb{R}^{d}\setminus N_{2}$ characterised the metric speed of a curve through the point $y$ depending on its derivative at $y$. In particular, we have shown that for almost every $y$, we have that if an absolutely continuous curve $\gamma:[0,1]\to\mathbb{R}^{d}$ passes through $y$ (say at time $t\in[0,1]$) and is differentiable at $t$, then the metric speed $|\dot{\gamma_{t}}|$ exists and is given by $|\dot{\gamma}_{t}|_{euc}$ if $y\in(0,1)^{d}$ and $x_{1}(y)\in O$, $\alpha(\dot{\gamma}_{t})$ if $y\in(0,1)^{d}$ and $x_{1}(y)\in\mathbb{R}\setminus(N^{(2)}\cup O)$, and $\sqrt{2}|\dot{\gamma}_{t}|_{euc}$ if $y\in([0,1]^{d})^{c}$.
Step 5. Example of functions that shows that the Cheeger energy does not satisfy the parallelogram identity.
This is the final step of the argument. Consider the smooth functions $f_{1},f_{2},f_{3},f_{4}:M=\mathbb{R}^{d}\to\mathbb{R}$ given by
@@
Now again pick a point $y\in(0,1)^{d}$ such that $x_{1}(y)$ is a Lebesgue point of $\{\theta=2\}$. Pick a vector $0\neq u\in\mathbb{R}^{d}$. Denote by $u_{2}$ the second component of $u$. Now, we may compute
@@
$\lim_{h\to 0}\frac{|f_{i}(\gamma^{y,u}_{h})-f_{i}(\gamma^{y,u}_{0})|}{\mathsf{d}_{g}(\gamma^{y,u}_{0},\gamma^{y,u}_{h})}=\frac{|\partial_{u}f_{i}|(\gamma_{0}^{y,u})}{\alpha(u)},\ \mathrm{for\ }1\leq i\leq 4.$
@@
$\frac{|\partial_{u}f_{1}|}{\alpha(u)}=\frac{|u_{1}|}{|u_{1}|+|u^{\prime}|_{euc}}\leq\frac{|u_{1}|}{2|u_{1}|+(|u^{\prime}|_{euc}-|u_{1}|)}\leq\frac{1}{2},$ (4.29)
@@
Here, in (4.29), we used that $|u_{1}|<|u^{\prime}|_{euc}$ which implies that $|u^{\prime}|_{euc}-|u_{1}|>0$. In (4.30), (4.31), and (4.32), one can see that $u_{(2)}:=e_{2}, u=u$$u_{(3)}:=e_{1}+e_{2}$, and $u=u_{(4)}:=e_{1}-e_{2}$ yield equality. We now define the functions $G_{1},G_{2},G_{3},G_{4}:\mathbb{R}^{d}\to[0,\infty]$ via
@@
Let $\gamma:[0,1]\to M$ be absolutely continuous and transversal to a $\mathcal{L}^{d}$-null set $N\supset N_{g}\cup N_{2}$. As $f_{i}\in C^{1}$, we may apply the chain rule. We have that for almost every $t\in[0,1]$, it holds that $\gamma_{t}\in([0,1]^{d})^{c}\cup(O\times(0,1)^{d-1})$ or $\gamma_{t}\in((0,1)\setminus O)\times(0,1)^{d-1}$ and $t$ is a differentiability point of $\gamma$. In the first case, we get that
$\Big|\frac{{\rm d}}{{\rm d}s}f_{i}\circ\gamma\Big|_{s=t}\Big|=|\partial_{\dot{\gamma}_{t}}f_{i}(\gamma_{t})|=|\langle\nabla_{g}f_{i}(\gamma_{t}),\dot{\gamma}_{t}\rangle_{g}|\leq|\nabla_{g}f_{i}|_{g}(\gamma_{t})|\dot{\gamma}_{t}|_{g}=G_{i}(\gamma_{t})|\dot{\gamma}_{t}|.$ (4.34)
@@
We next want to see that these are indeed the minimal weak upper gradients. For that we need the following result from [30, Lemma 4.23].
@@
$F_{v,\delta,x}:t\to\frac{1}{\mathcal{L}^{d}(B_{\delta}(x))}\int_{B_{\delta}(x)}k(y+tv)\,{\rm d}\mathcal{L}^{d}(y),\quad\forall t,\;|t|<1.$
@@
$F_{u_{(i)},\delta,x}:t\to\frac{1}{\mathcal{L}^{d}(B_{\delta}(x))}\int_{B_{\delta}(x)}|Df_{i}|_{w}(y+tu_{(i)})\,{\rm d}\mathcal{L}^{d}(y),\ \mathrm{and}$
$G_{u_{(i)},\delta,x}:t\to\frac{1}{\mathcal{L}^{d}(B_{\delta}(x))}\int_{B_{\delta}(x)}G_{i}(y+tu_{(i)})\,{\rm d}\mathcal{L}^{d}(y),$
@@
$\frac{1}{\mathcal{L}^{d}(B_{\delta}(x))}\int_{B_{\delta}(x)}||Df_{i}|_{w}(y)-|Df_{i}|_{w}(x)|\,{\rm d}\mathcal{L}^{d}(y)\leq\frac{\varepsilon}{8},\ \mathrm{and}$ $\frac{1}{\mathcal{L}^{d}(B_{\delta}(x))}\int_{B_{\delta}(x)}|G_{i}(y)-G_{i}(x)|\,{\rm d}\mathcal{L}^{d}(y)\leq\frac{\varepsilon}{8}.$ (4.37)
@@
$\frac{1}{2\tau}\frac{1}{\mathcal{L}^{d}(B_{\delta}(x))}\int_{-\tau}^{\tau}\Bigg|\int_{B_{\delta}(x)}|Df_{i}|_{w}(y+tu_{(i)})\,{\rm d}\mathcal{L}^{d}(y)-\int_{B_{\delta}(x)}|Df_{i}|_{w}(y)\,{\rm d}\mathcal{L}^{d}(y)\Bigg|\,{\rm d}t\leq\frac{\varepsilon}{8},\ \mathrm{and}$
$\frac{1}{2\tau}\frac{1}{\mathcal{L}^{d}(B_{\delta}(x))}\int_{-\tau}^{\tau}\Bigg|\int_{B_{\delta}(x)}G_{i}(y+tu_{(i)})\,{\rm d}\mathcal{L}^{d}(y)-\int_{B_{\delta}(x)}G_{i}(y)\,{\rm d}\mathcal{L}^{d}(y)\Bigg|\,{\rm d}t\leq\frac{\varepsilon}{8}.$ (4.38)
@@
$\frac{1}{2\tau}\frac{1}{\mathcal{L}^{d}(B_{\delta}(x))}\int_{-\tau}^{\tau}\int_{B_{\delta}(x)}G_{i}(y+tu_{(i)})-|Df_{i}|_{w}(y+tu_{(i)})\,{\rm d}\mathcal{L}^{d}(y){\rm d}t\geq\varepsilon-\frac{4\varepsilon}{8}=\frac{\varepsilon}{2}.$ (4.39)
@@
$=\frac{1}{\mathcal{L}^{d}(B_{\delta}(x))}\int_{-\tau}^{\tau}\int_{B_{\delta}(x)}(G_{i}(y+tu_{(i)})-|Df_{i}|_{w}(y+tu_{(i)}))|\dot{\gamma}^{y,i}_{t}|\,{\rm d}\mathcal{L}^{d}(y){\rm d}t\geq\frac{\tau\varepsilon}{\sqrt{2}}.$ (4.43)
Now we observe that for all points $y\in\mathbb{R}^{d}$, $t\in[0,1]$ such that $z:=y+tu_{(i)}\in(O\times(0,1)^{d-1})\cup([0,1]^{d})^{c}$, we have that $g\equiv\mathrm{Id}$ a neighbourhood of $z$ and $\nabla_{g}f_{i}(z)=u_{(i)}$. Then by definition we have that $G_{i}(z)|\dot{\gamma}_{t}^{y,i}|=G_{i}|u_{(i)}|_{euc}=\langle\nabla_{euc}f_{i},u_{(i)}\rangle_{euc}=\partial_{u_{(i)}}f_{i}.$ Similarly, we have that for all points $y\in\mathbb{R}^{d}$, $t\in[0,1]$ such that $z:=y+tu_{(i)}\in(N_{2}\cup N_{g})^{c}\setminus(O\times(0,1)^{d-1}\cup([0,1]^{d})^{c})$ that $\partial_{u_{(i)}}f_{i}(z)=G_{i}(z)|\dot{\gamma}_{t}^{y,i}|$ (see (4.35) and the equality cases of (4.25), (4.30), (4.31), and (4.32)).
@@
Pick a function $\phi\in C_{c}^{\infty}(\mathbb{R}^{d})$ such that $\phi\equiv 1$ on $1,2]^{d}$. Consider the functions $\tilde{f}_{i}:=\phi f_{i}$ for $1\leq i\leq 4$. Then by the previous observations, we have that
@@
Moreover, we have that $\tilde{f}_{3}=\tilde{f}_{1}+\tilde{f}_{2}$ and $\tilde{f}_{4}=\tilde{f}_{1}-\tilde{f}_{2}$ and all four functions have finite Cheeger energy. But as $\mathcal{L}^{d}((0,1)^{d}\setminus O\times(0,1)^{d-1})\geq 1-2\kappa>0$ and the fact that $\mathrm{vol}_{g}$ and $\mathcal{L}^{d}$ are $2^{\frac{d}{2}}$-equivalent, we get that
@@ 4.3 An example of a Sobolev Riemannian metric, neither quasi-Riemannian nor infinitesimally Hilbertian
Fix a $d\geq 3$ and set p:=d1p:=d-1. We will now refine the previous examples to construct a Riemannian metric $g$ on $\mathbb{R}^{d}$ such that $g,g^{-1}\in L^{\infty}$, $g\in W^{1,p}_{\rm loc}$ such that $(M,\mathsf{d}_{g},g)$ is not quasi-Riemannian and such that the metric measure space $(M,\mathsf{d}_{g},\mathrm{vol}_{g})$ is not infinitesimally Hilbertian.
Define $\omega_{d-1}:=\mathcal{L}^{d-1}(B^{\mathbb{R}^{d-1}}_{1}(0))$. We start with the following preparatory result:
Lemma 4.7.
For each $R>0$ and $\varepsilon>0$, there exists an $r>0$ and a function $f\in C^{0,1}_{c}(B^{\mathbb{R}^{d-1}}_{R}(0),[0,1])$ with $f(B_{r}(0))=\{1\}$ and such that
$\int_{\mathbb{R}^{d-1}}|\nabla_{euc}f|^{d-1}\,{\rm d}x<\varepsilon.$ (4.49)
Proof.
This proof uses the idea of the proof of [13, Theorem 4.16]. For each $\rho>0$ consider the function $\eta_{\rho}\in C_{c}^{0,1}(B^{\mathbb{R}^{d-1}}_{\rho}(0),[0,1])$ defined by
$\eta_{\rho}(y):=\left\{\begin{array}[]{ll}1&\mathrm{if}\ y\in B_{\rho/4}(0),\\ 1-\frac{4(|y|-\frac{\rho}{4})}{\rho}&\mathrm{if}\ y\in B_{\rho/2}(0)\setminus B_{\rho/4}(0)\\ 0&\mathrm{if}\ y\in B^{c}_{\rho/2}(0).\end{array}\right$. (4.53)
We can estimate that
$\int_{\mathbb{R}^{d-1}}|\nabla_{euc}\eta_{\rho}|^{d-1}\,{\rm d}x\leq\int_{B_{\rho/2}(0)}\frac{4^{d-1}}{\rho^{d-1}}\leq 2^{d-1}\omega_{d-1}.$ (4.54)
Define
$S_{j}:=\sum_{k=1}^{j}\frac{1}{k}.$
Now for each $j\in\mathbb{N}$ define the function $f_{j}\in C^{0,1}_{c}(B_{R}(0),[0,1])$ as
$f_{j}:=\frac{1}{S_{j}}\sum_{k=1}^{j}\frac{\eta_{R/8^{k}}}{k}.$ (4.55)
Then the construction of $\eta_{\rho}$ implies that on $B_{R/8^{j+1}}(0)$ we have that $f_{j}\equiv 1$. Moreover, we have that $\mathrm{supp}\,\nabla_{euc}\eta_{R/8^{k}}=\overline{B_{R/(2\cdot 8^{k})}(0)}\setminus B_{R/(4\cdot 8^{k})}(0)$, hence for any $k\neq k^{\prime}\in\mathbb{N}$, we get that $\mathrm{supp}\,\nabla_{euc}\eta_{R/8^{k}}\cap\mathrm{supp}\,\nabla_{euc}\eta_{R/8^{k^{\prime}}}=\emptyset$. We may now estimate
$\int_{\mathbb{R}^{d-1}}|\nabla_{euc}f_{j}|^{d-1}\,{\rm d}x=\frac{1}{S_{j}^{d-1}}\sum_{k=1}^{j}\frac{1}{k^{d-1}}\int_{\mathbb{R}^{d-1}}|\nabla_{euc}\eta_{R/8^{k}}|^{d-1}\,{\rm d}x\leq\frac{2^{d-1}\omega_{d-1}}{S_{j}^{d-1}}\sum_{k=1}^{j}\frac{1}{k^{d-1}}.$
Now as d1>1d-1>1, we get that $\lim_{j\to\infty}\frac{2^{d-1}\omega_{d-1}}{S_{j}^{d-1}}\sum_{k=1}^{j}\frac{1}{k^{d-1}}=0$. Hence, we can find a $j\in\mathbb{N}$ such that $\int_{\mathbb{R}^{d-1}}|Df_{j}|^{d-1}\,{\rm d}x<\varepsilon$. Then setting $r:=\frac{R}{8^{j+1}}$ finishes the proof. ∎
Pick a decreasing sequence $(R_{m})_{m\in\mathbb{N}}\subset(0,1)$ such that $R_{0}=\frac{1}{2}$ and for $m\geq 1$ it holds
@@
We inductively define the finite sets $D_{m}\subset(0,1)^{d-1}$, open sets $O_{m}\subset(-1,2)^{d-1}$, a decreasing null sequence $(r_{m})_{m\in\mathbb{N}}\subset(0,1)$ and functions $\theta_{m}\in C^{\infty}(\mathbb{R}^{d-1})$ as follows:
$D_{0}:=O_{0}:=\emptyset,\ \theta_{0}\equiv 2\in C^{\infty}(\mathbb{R}^{d-1}),r_{0}:=\frac{1}{4}.$
@@
Moreover, note that from $D_{m}\subset\frac{1}{2^{m+1}}\mathbb{Z}^{d-1}$, together with (4.56) and (4.59), we have that for $^{\prime}\in D_{m}, xxx\neq x^{\prime}$, it holds
@@
Using Lemma 4.7, we can for each $m\in\mathbb{N}$ find a function $\eta_{m}\in C_{c}^{0,1}(\mathbb{R}^{d-1},[0,1])$ such that there exists an $\rho_{m}\in(0,r_{m}/5)$ with $\eta_{m}|_{B_{\rho_{m}}(0)}\equiv 1$, $\eta_{m}|_{\mathbb{R}^{d-1}\setminus B_{r_{m}/5}(0)}\equiv 2$ and
$2^{(m+1)(d+1)}\int_{\mathbb{R}^{d-1}}|D\eta_{m}|^{d-1}\,{\rm d}x\leq 2^{-m}.$ (4.63)
@@
$\psi_{m}(y):=\left\{\begin{array}[]{ll}2&\mathrm{if}\ y\in O_{m}^{c},\\ \eta_{m}\big(y-x\big)&\mathrm{if}\ y\in B_{r_{m}}(x)\ \mathrm{for\ some\ }x\in D_{m}.\end{array}\right$. (4.66)
@@
$\lim_{r\to 0}\frac{\sum_{m\in\mathbb{N}}r_{m}\#\{x\in D_{m}:\overline{B_{r_{m}/5}(x)}\cap\overline{B_{r}(y)}\neq\emptyset\}}{r}=0\ \mathrm{and}$ (4.72)
$\lim_{r\to 0}\frac{\sum_{m\in\mathbb{N}}r^{d-1}_{m}\#\{x\in D_{m}:\overline{B_{r_{m}/5}(x)}\cap\overline{B_{r}(y)}\neq\emptyset\}}{r^{d-1}}=0.$ (4.73)
@@
$\zeta_{m}(y,r):=\#\{x\in D_{m}:\overline{B_{r_{m}/5}(x)}\cap\overline{B_{r}(y)}\neq\emptyset\}.$ (4.74)
@@
$\zeta_{m}(y,r)$ $\leq\#\big\{\frac{1}{2^{m+1}}\mathbb{Z}^{d-1}\cap\overline{B_{r+\frac{r_{m}}{5}}(y)}\neq\emptyset\big\}\leq C(d)2^{(d-1)(m+1)}\frac{5^{d-1}}{4^{d-1}}r^{d-1},$
@@
$\frac{\sum_{m\in\mathbb{N}}r^{d-1}_{m}\#\{x\in D_{m}:\overline{B_{r_{m}/5}(x)}\cap\overline{B_{r}(y)}\neq\emptyset\}}{r^{d-1}}$ $\leq C(d)\sum_{m:\,4r_{m}\leq 5r}2^{(d-1)(m+1)}r_{m}^{d-1}$
$\leq C(d)\sum_{m:\,4r_{m}\leq 5r}2^{(d-1)(m+1)}r_{m}.$ (4.76)
Now (4.57) implies that the right-hand-side of (4.76) converges to 0 as $r\to 0$, which implies (4.73). Moreover,
$\frac{\sum_{m\in\mathbb{N}}r_{m}\#\{x\in D_{m}:\overline{B_{r_{m}/5}(x)}\cap\overline{B_{r}(y)}\neq\emptyset\}}{r}$ $\leq r^{d-2}C(d)\sum_{m:\,4r_{m}\leq 5r}2^{(d-1)(m+1)}r_{m},$ (4.77)